OpenStudy (anonymous):

jacob travels for 5 hours at a constant speed.The return trip takes 4 hours, traveling 15 mph faster,also at a constant speed.What was the distance each way and what were his speeds for each way?

OpenStudy (anonymous):

use this formula d=rt

OpenStudy (anonymous):

so 5x+4x+15=d

OpenStudy (anonymous):

v2=v1+15 5v1=4(v1+15) 300 miles 60 MPH 75 MPH

OpenStudy (anonymous):

please elaborate nsutton can i use x instead of v

OpenStudy (anonymous):

I like to use v for velocity and x as a displacement, but that is only personal preference. the method is the same.

OpenStudy (anonymous):

please break it down more

OpenStudy (anonymous):

v2 is return velocity, v1 is first velocity. solve for v1. solve for v2 using the value of v1. v1*5 gives the distance.

OpenStudy (anonymous):

so if v1 is first velosity why wouldnt it be v1-15

OpenStudy (anonymous):

v2 is v1+15mph

OpenStudy (anonymous):

speed=distance/time. let v1 be his first velocity. x be the distance travelled in the two cases. t be the time taken. v1=x/5 in first case. x=5*v........................(1) in second case v2=v1+15. v1+15=x/4 x=(v1+15)*4...............(2). equate 1 and 2. on simplification u get v1=15km/hr. v2=15+15=30km/hr. substitute v1 value in(1). u get x=75km.

OpenStudy (anonymous):

in (1) x=v1*5.

OpenStudy (anonymous):

Re-simplify your equation @anusha.p

OpenStudy (anonymous):

you did not distribute the four to the fifteen before grouping like terms.

OpenStudy (anonymous):

yup.v1=60. v2=75. x=300. i'm srry guys.

OpenStudy (anonymous):

it's all good.

OpenStudy (anonymous):

ty nsutton and anusha.p

OpenStudy (anonymous):

and it is not km/hr.infact miles/hr.